Please help:
Five silver coins and eight gold coins are in a purse. Four coins are to be selected at random, without replacement. Determine the probability that two silver and two gold coins are selected.

Answer by rothauserc(4718)

there are 13 coins in the purse
:
the probability of selecting two silver coins is 5/13 * 4/12 = 0.1282
the probability of selecting two gold coins is 8/11 * 7/10 = 0.5091
:
the probability of selecting two silver and two gold coins without replacement is 0.1282 * 0.5091 = 0.0653
: